One hundred liters of a salt and water solution contains 1% salt. After some of the water has evaporated, the solution contains 5% salt. How many liters of water evaporated?
Answer by TimothyLamb(4379)

x = liters of water evaporated
---
quantity of salt equation:
---
0.01*100 = 0.05*(100 - x)
1 = 5 - 0.05x
0.05x = 4
x = 4/0.05
x = 80
---
Free algebra tutoring live chat:
https://sooeet.com/chat.php?gn=algebra
---
Solve and graph linear equations:
https://sooeet.com/math/linear-equation-solver.php
---
Solve quadratic equations with quadratic formula:
https://sooeet.com/math/quadratic-formula-solver.php
---
Solve systems of linear equations up to 6-equations 6-variables:
https://sooeet.com/math/system-of-linear-equations-solver.php
---